# Bonsai: Gradient-free Graph Distillation for Node Classification

### v1

@@ 4.3 Cross-architecture Generalization
Dataset Gnn Random Herding GCond GDEM GCSR Bonsai Full
Gat $42.73{\scriptstyle\pm 1.03}$ $46.09{\scriptstyle\pm 0.86}$ $35.24{\scriptstyle\pm 0.00}$ $73.49{\scriptstyle\pm 2.64}$ $37.60{\scriptstyle\pm 1.34}$ $78.67{\scriptstyle\pm 0.89}$ $85.70{\scriptstyle\pm 0.09}$
Cora Gin $50.48{\scriptstyle\pm 0.85}$ $33.80{\scriptstyle\pm 2.42}$ $33.91{\scriptstyle\pm 1.23}$ $75.92{\scriptstyle\pm 4.24}$ $60.70{\scriptstyle\pm 4.44}$ $84.80{\scriptstyle\pm 0.41}$ $86.62{\scriptstyle\pm 0.28}$
Gat $46.19{\scriptstyle\pm 1.38}$ $52.07{\scriptstyle\pm 0.11}$ $21.47{\scriptstyle\pm 0.00}$ $23.87{\scriptstyle\pm 2.76}$ $21.50{\scriptstyle\pm 0.06}$ $69.43{\scriptstyle\pm 0.82}$ $77.48{\scriptstyle\pm 0.75}$
CiteSeer Gin $47.90{\scriptstyle\pm 0.65}$ $39.67{\scriptstyle\pm 0.82}$ $19.49{\scriptstyle\pm 1.09}$ $67.64{\scriptstyle\pm 4.45}$ $64.74{\scriptstyle\pm 1.88}$ $72.16{\scriptstyle\pm 0.60}$ $75.45{\scriptstyle\pm 0.23}$
Gat $78.85{\scriptstyle\pm 0.09}$ $76.64{\scriptstyle\pm 0.02}$ $41.55{\scriptstyle\pm 3.18}$ $80.75{\scriptstyle\pm 0.47}$ $40.47{\scriptstyle\pm 0.00}$ $85.88{\scriptstyle\pm 0.28}$ $86.33{\scriptstyle\pm 0.08}$
Pubmed Gin $78.43{\scriptstyle\pm 0.22}$ $62.22{\scriptstyle\pm 0.13}$ $32.84{\scriptstyle\pm 6.27}$ $78.72{\scriptstyle\pm 0.95}$ $33.75{\scriptstyle\pm 5.58}$ $85.57{\scriptstyle\pm 0.26}$ $84.66{\scriptstyle\pm 0.05}$
Gat $43.56{\scriptstyle\pm 1.06}$ $36.34{\scriptstyle\pm 1.14}$ $40.85{\scriptstyle\pm 1.08}$ $18.44{\scriptstyle\pm 9.42}$ OOT $45.62{\scriptstyle\pm 1.85}$ $51.42{\scriptstyle\pm 0.07}$
Flickr Gin $42.90{\scriptstyle\pm 0.76}$ $41.87{\scriptstyle\pm 4.52}$ $16.65{\scriptstyle\pm 6.55}$ $19.44{\scriptstyle\pm 9.68}$ OOT $44.90{\scriptstyle\pm 0.88}$ $45.37{\scriptstyle\pm 0.57}$
Figure 4: (a) Distillation times of various baselines on Gcn at a compression ratio of $0.5\%$, (b) Impact of the various components within Bonsai on its performance.
4.4 Distillation and Training Efficiency
Distillation efficiency: Fig. LABEL:fig:time presents the time consumed by Bonsai and other baselines. We note that while Bonsai is CPU-bound, all other algorithms are GPU-bound. Despite being CPU-bound, Bonsai, on average, is more than $22$-times faster than the closest baseline Gdem. Gdem has a quadratic time complexity due to being reliant on eigen decomposition. In contrast, Bonsai is linear to the number of nodes and edges in the graph. In addition, all three baselines require training on the full dataset. Gdem incurs a smaller cost by training an MLP solely on node attributes to predict node labels as a pre-processing step. However, GCond and Gcsr perform full Gnn training on the entire training set. This approach essentially negates the purpose of graph distillation, as it involves the very computational burden we aim to avoid. GCond is marginally faster than full training due to employing some Gcn-specific optimizations that outperform the native Gcn implementation available in PyG v2.3.1, which we use for full dataset training. Due to the same reason, Gcsr, although performing full-dataset training 100 times, does not incur a 100-fold increase in running time.
Training efficiency: Fig. 5 presents the training times on the three largest distilled datasets in Flickr, Ogbn-arxiv and Flickr. Results on other datasets are presented in Figs. 8, 9 and 10 in the App. Two key patterns emerge. Bonsai is faster across all architectures and datasets. Second, the gap between Bonsai and the baselines increase with increase in distilled dataset size. This is explained from the property that all baselines generate a fully connected weighted graph. Hence, the $L$-hop neighborhood of a node is significantly larger leading to higher computation and storage requirements. In contrast, through PPR, Bonsai sparsifies the graph by pruning the nodes with small impacts on embeddings of exemplars.
@@ 4.4 Distillation and Training Efficiency
Figure 5: Training times on the distilled datasets produced by Bonsai and the considered baselines on the three largest datasets at $0.5\%$ compression. Note that there is no bar for Gcsr in Ogbn-arxiv and Reddit since Gcsr fails to create a distilled dataset within 5 hours.
@@ 4.5 Ablation Study and Impact of Parameters
Fig. LABEL:fig:ablation presents an ablation study comparing the full Bonsai method against two variants: Bonsai-PPR, which omits the use of Personalized PageRank, and Bonsai-Rev-$k$-NN, which substitutes the Rev-$k$-NN-based coverage maximization with random exemplar selection. The results reveal that the Rev-$k$-NN component generally has a more substantial impact on performance, with the exception of the Reddit dataset where PPR marginally dominates.
This observation aligns with expectations, as random exemplar selection would fail to comprehensively represent the input space. The PPR component, while important, typically has a lesser impact since its primary function is to sparsify the initially distilled graph by eliminating message passing paths that minimally affect the embeddings of selected exemplars.
The Reddit dataset presents a unique case where PPR has a more significant impact. This is attributed to the dataset’s high average node degree of 100100~{}100100, in contrast to less than $14$ in other networks. Consequently, within the given budget, very few exemplars can be selected due to their substantially larger $L$-hop neighborhoods. In this scenario, PPR plays a crucial role by pruning these expansive neighborhoods, thereby creating space for additional exemplars and improving overall performance.
@@ 5 Conclusions and Future Works
In this work, we have developed Bonsai, a novel graph distillation method that addresses critical limitations in existing approaches. By leveraging the fundamental role of computation trees in message-passing Gnns, Bonsai achieves superior performance in node classification tasks across multiple real-world datasets. Our method stands out as the first linear-time, model-agnostic graph distillation algorithm, offering significant improvements in both accuracy and computational efficiency. Bonsai’s unique approach of encoding exemplar trees that maximize representation of the full training set’s computation trees has proven to be highly effective. This strategy not only overcomes the paradoxical requirement of training on full datasets for distillation but also eliminates the need for repeated distillation when changing hyper-parameters or Gnn architectures. Furthermore, Bonsai achieves substantial size reduction without resorting to fully-connected, edge-weighted graphs, thereby reducing computational demands. In contrast to baselines, Bonsai is completely CPU-bound, resulting in a smaller greenhouse footprint along with reduced computation demands, making it a more environmentally friendly option. These features collectively position Bonsai as a significant advancement in the field of graph distillation, offering both performance benefits and sustainability advantages.
Limitations and Future Works: While existing research on graph distillation has primarily focused on node and graph classification tasks, Gnns have demonstrated their versatility across a broader spectrum of applications. In our future work, we aim to expand the scope of graph distillation by developing task-agnostic data distillation algorithms.
@@
To support the reproducibility of our work, we provide several resources in the paper and its supplementary materials. The source code of our models and algorithms is available at https://anonymous.4open.science/r/bonsai, which also includes details of how to train the baseline models in different settings. All theoretical results and assumptions are detailed in the Appendix A, ensuring clarity in our claims. Full experimental settings are documented in Appendix B.
@@ A.3 Proofs of Monotonicity, Submodularity
A.4 Complexity Analysis
@@
    Coverage Maximization: Each iteration in Alg. 1, iterates over all nodes (trees) in the graph ($\mathbb{T}$) that have not yet been added to the set of exemplars. For each node (equivalently, tree rooted at this node), we compute its marginal gain (line 3 in Alg. 1), which is bounded by the sample size $z$, since the cardinality of a Rev-$k$-NN set is bounded by the number of trees sampled for Rev-$k$-NN computation ((§ 3.1.1)). Since $z\ll n$, the complexity per iteration is $\mathcal{O}(zn)\approx\mathcal{O}(n)$. The number of iterations is bounded by the size of the distillation budget, which we typically set to less than $3\%$ of the full dataset and hence has negligible impact on the complexity.
@@
    PPR: PPR can be computed in linear time (Lofgren et al., 2014). Since the size of the distilled set is at most the full graph, this bounds the cost of PPT to $\mathcal{O}(n)$.

### v3

@@ 4.3 Cross-architecture Generalization
Dataset % Gnn Random Herding GCond GDEM GCSR Bonsai Full
0.5 Gat $41.44{\scriptstyle\pm 1.73}$ $33.80{\scriptstyle\pm 0.07}$ $13.21{\scriptstyle\pm 1.99}$ $63.91{\scriptstyle\pm 5.91}$ $15.09{\scriptstyle\pm 6.19}$ $75.42{\scriptstyle\pm 1.61}$
1 Gat $42.73{\scriptstyle\pm 1.03}$ $46.09{\scriptstyle\pm 0.86}$ $35.24{\scriptstyle\pm 0.00}$ $73.49{\scriptstyle\pm 2.64}$ $37.60{\scriptstyle\pm 1.34}$ $78.67{\scriptstyle\pm 0.89}$ $85.70{\scriptstyle\pm 0.09}$
Cora 3 Gat $60.22{\scriptstyle\pm 0.67}$ $56.75{\scriptstyle\pm 0.45}$ $35.24{\scriptstyle\pm 0.00}$ $75.28{\scriptstyle\pm 4.86}$ $36.72{\scriptstyle\pm 0.81}$ $80.66{\scriptstyle\pm 0.80}$
0.5 Gin $49.04{\scriptstyle\pm 0.50}$ $34.39{\scriptstyle\pm 1.03}$ $14.13{\scriptstyle\pm 6.80}$ $63.65{\scriptstyle\pm 7.11}$ $76.05{\scriptstyle\pm 0.44}$ $85.42{\scriptstyle\pm 0.74}$
1 Gin $50.48{\scriptstyle\pm 0.85}$ $33.80{\scriptstyle\pm 2.42}$ $33.91{\scriptstyle\pm 1.23}$ $75.92{\scriptstyle\pm 4.24}$ $60.70{\scriptstyle\pm 4.44}$ $84.80{\scriptstyle\pm 0.41}$ $86.62{\scriptstyle\pm 0.28}$
3 Gin $59.52{\scriptstyle\pm 0.88}$ $36.35{\scriptstyle\pm 0.59}$ $31.70{\scriptstyle\pm 4.97}$ $59.59{\scriptstyle\pm 7.95}$ $51.62{\scriptstyle\pm 5.00}$ $85.42{\scriptstyle\pm 0.53}$
0.5 Gat $42.76{\scriptstyle\pm 0.35}$ $36.04{\scriptstyle\pm 0.46}$ $21.47{\scriptstyle\pm 0.00}$ $69.86{\scriptstyle\pm 2.28}$ $21.92{\scriptstyle\pm 0.76}$ $68.56{\scriptstyle\pm 0.57}$
1 Gat $46.19{\scriptstyle\pm 1.38}$ $52.07{\scriptstyle\pm 0.11}$ $21.47{\scriptstyle\pm 0.00}$ $23.87{\scriptstyle\pm 3.05}$ $21.50{\scriptstyle\pm 0.06}$ $69.43{\scriptstyle\pm 0.82}$ $77.48{\scriptstyle\pm 0.75}$
CiteSeer 3 Gat $61.65{\scriptstyle\pm 0.51}$ $65.17{\scriptstyle\pm 0.00}$ $21.26{\scriptstyle\pm 0.22}$ $22.90{\scriptstyle\pm 1.20}$ $21.50{\scriptstyle\pm 0.06}$ $69.94{\scriptstyle\pm 1.15}$
0.5 Gin $44.86{\scriptstyle\pm 0.43}$ $22.97{\scriptstyle\pm 0.30}$ $21.47{\scriptstyle\pm 0.00}$ $67.69{\scriptstyle\pm 3.28}$ $50.66{\scriptstyle\pm 1.17}$ $71.80{\scriptstyle\pm 0.26}$
1 Gin $47.90{\scriptstyle\pm 0.65}$ $39.67{\scriptstyle\pm 0.82}$ $19.49{\scriptstyle\pm 1.09}$ $67.64{\scriptstyle\pm 4.45}$ $64.74{\scriptstyle\pm 1.88}$ $72.16{\scriptstyle\pm 0.60}$ $75.45{\scriptstyle\pm 0.23}$
3 Gin $61.83{\scriptstyle\pm 0.68}$ $60.48{\scriptstyle\pm 0.26}$ $18.65{\scriptstyle\pm 2.56}$ $48.65{\scriptstyle\pm 8.17}$ $59.95{\scriptstyle\pm 9.07}$ $70.51{\scriptstyle\pm 0.54}$
0.5 Gat $77.73{\scriptstyle\pm 0.12}$ $75.44{\scriptstyle\pm 0.02}$ $37.49{\scriptstyle\pm 4.01}$ $80.06{\scriptstyle\pm 1.16}$ $38.29{\scriptstyle\pm 8.13}$ $85.66{\scriptstyle\pm 0.38}$
1 Gat $78.85{\scriptstyle\pm 0.09}$ $76.64{\scriptstyle\pm 0.02}$ $41.55{\scriptstyle\pm 3.18}$ $80.75{\scriptstyle\pm 0.47}$ $40.47{\scriptstyle\pm 0.00}$ $85.88{\scriptstyle\pm 0.28}$ $86.33{\scriptstyle\pm 0.08}$
PubMed 3 Gat $82.84{\scriptstyle\pm 0.11}$ $78.48{\scriptstyle\pm 0.03}$ $37.77{\scriptstyle\pm 3.61}$ $65.08{\scriptstyle\pm 9.53}$ $40.27{\scriptstyle\pm 0.20}$ $85.62{\scriptstyle\pm 0.36}$
0.5 Gin $77.45{\scriptstyle\pm 0.14}$ $48.48{\scriptstyle\pm 1.33}$ $30.91{\scriptstyle\pm 4.57}$ $78.78{\scriptstyle\pm 0.91}$ $36.88{\scriptstyle\pm 12.06}$ $84.32{\scriptstyle\pm 0.33}$
1 Gin $78.43{\scriptstyle\pm 0.22}$ $62.22{\scriptstyle\pm 0.13}$ $32.84{\scriptstyle\pm 6.27}$ $78.72{\scriptstyle\pm 0.95}$ $33.75{\scriptstyle\pm 5.58}$ $85.57{\scriptstyle\pm 0.26}$ $84.66{\scriptstyle\pm 0.05}$
3 Gin $80.56{\scriptstyle\pm 0.17}$ $45.40{\scriptstyle\pm 0.46}$ $36.11{\scriptstyle\pm 3.47}$ $81.08{\scriptstyle\pm 0.99}$ $32.01{\scriptstyle\pm 6.77}$ $85.66{\scriptstyle\pm 0.23}$
0.5 Gat $43.64{\scriptstyle\pm 0.99}$ $36.50{\scriptstyle\pm 13.22}$ $40.24{\scriptstyle\pm 3.20}$ $25.43{\scriptstyle\pm 10.37}$ $28.03{\scriptstyle\pm 6.60}$ $48.22{\scriptstyle\pm 3.60}$
1 Gat $43.56{\scriptstyle\pm 1.06}$ $36.34{\scriptstyle\pm 1.14}$ $40.85{\scriptstyle\pm 1.08}$ $18.44{\scriptstyle\pm 9.42}$ OOT $45.62{\scriptstyle\pm 1.85}$ $51.42{\scriptstyle\pm 0.07}$
Flickr 3 Gat $45.71{\scriptstyle\pm 1.87}$ $42.70{\scriptstyle\pm 1.17}$ $41.51{\scriptstyle\pm 9.81}$ $25.83{\scriptstyle\pm 11.39}$ OOT $47.80{\scriptstyle\pm 2.06}$
0.5 Gin $42.67{\scriptstyle\pm 0.83}$ $39.98{\scriptstyle\pm 7.21}$ $13.65{\scriptstyle\pm 7.54}$ $14.10{\scriptstyle\pm 5.68}$ $05.92{\scriptstyle\pm 1.01}$ $44.97{\scriptstyle\pm 2.23}$
1 Gin $42.90{\scriptstyle\pm 0.76}$ $41.87{\scriptstyle\pm 4.52}$ $16.65{\scriptstyle\pm 6.55}$ $19.44{\scriptstyle\pm 9.68}$ OOT $44.90{\scriptstyle\pm 0.88}$ $45.37{\scriptstyle\pm 0.57}$
3 Gin $19.63{\scriptstyle\pm 4.21}$ $43.72{\scriptstyle\pm 3.26}$ $24.25{\scriptstyle\pm 14.43}$ $20.97{\scriptstyle\pm 6.64}$ OOT $45.04{\scriptstyle\pm 1.94}$
4.4 Condensation Efficiency
Table. 6 presents the time consumed by Bonsai and other baselines. We note that while Bonsai is CPU-bound, all other algorithms are GPU-bound. Despite being CPU-bound, Bonsai, on average, is more than $7$-times faster than the fastest baseline Exgc. In addition, all baselines require training on the full dataset. This algorithm design shows up in the running times where the condensation time is higher than the full dataset training time, negating the very purpose of graph condensation. More worryingly, the condensation process of all algorithms, except Bonsai, have higher carbon emissions than training on the full dataset (See Table 8 in Appendix).
Table 6: Condensation times of various methods in seconds at 0.5%. Condensation times that are higher than training on the full dataset itself are highlighted in red. The fastest condensation time for each dataset is highlighted in green.
Dataset GCond Gdem Gcsr Exgc GC-Sntk Geom Bonsai Full
Cora 2738 105 5260 34.87 82 12996 2.60 24.97
Citeseer 2712 167 6636 34.51 124 15763 2.75 24.87
Pubmed 2567 530 1319 114.96 117 OOT 24.24 51.06
Flickr 1935 3405 17445 243.28 612 OOT 118.23 180.08
Ogbn-arxiv 14474 569 OOT 1594.83 12218 OOT 298.64 524.67
Reddit 30112 20098 OOT 6903.47 29211 OOT 1170.64 2425.68
@@ 4.5 Ablation Study and Impact of Parameters
Fig. 4 presents an ablation study comparing Bonsai against three variants: (i) Bonsai-PPR, which omits the use of Personalized PageRank, (ii) Bonsai-Rev-$k$-NN, which substitutes the Rev-$k$-NN-based coverage maximization with random exemplar selection, but performs PPR, (iii) and random that does not do either PPR or Rev-$k$-NN. Before analyzing the trends, we emphasize an important distinction between Bonsai-Rev-$k$-NN and Random. While in Bonsai-Rev-$k$-NN, we randomly add computation trees till the condensation budget is exhausted, in Random, we iteratively add random nodes, till their induced subgraph exhausts the budget. Consequently, Random covers more diverse nodes with partial local neighborhood information, while Bonsai-Rev-$k$-NN selects a smaller node set with complete $L$-hop topology, enabling precise Gnn embedding computation.
Several important insights emerge from this experiment. Random performs significantly worse than Bonsai, demonstrating that the condensed graph’s information content substantially exceeds that of an equally sized subgraph induced by random node selection. This finding, consistent with the trend observed in Table 4, underscores the effectiveness of Bonsai over random sampling.
Second, removal of either Rev-$k$-NN or PPR result in significant drops across most datasets, indicating both play important and complementary roles. While Rev-$k$-NN identifies the exemplars that are important, PPR sparsifies the graph making space for more exemplars to be added. Interestingly, the relative impact of these components varies with dataset compression and size. At higher compression rates (such as 0.5%), removing Rev-$k$-NN leads to higher drop in accuracy across most datasets. However, this trend reverses as the condensed dataset size increases. Closer examination reveals that higher-ranked exemplars typically have higher degrees in their $L$-hop neighborhoods. When the memory budget is constrained, high-ranked exemplars are often skipped because their $L$-hop neighborhoods exceed the available memory. This limitation diminishes as the budget increases. Consequently, at higher budgets, edge density increases, making graph sparsification (via PPR) increasingly critical. This explains why PPR becomes more important at higher budgets or in inherently dense graphs like Reddit.
Finally, we discuss the performance comparison of Random with Bonsai-Rev-$k$-NN. This is an interesting comparison since both represent two distinct mechanisms of random selection. In most cases, Bonsai-Rev-$k$-NN achieves a higher accuracy indicating that obtaining full $L$-hop topological information for a smaller set of nodes leads to better results than partial $L$-hop information over a broader set of nodes.
Figure 4: Ablation study of Bonsai.
@@ 5 Conclusions and Future Works
In this work, we have developed Bonsai, a novel graph condensation method that addresses critical limitations in existing approaches. By leveraging the fundamental role of computation trees in message-passing Gnns, Bonsai achieves superior performance in node classification tasks across multiple real-world datasets. Our method stands out as the first linear-time, model-agnostic graph condensation algorithm, offering significant improvements in both accuracy and computational efficiency. Bonsai’s unique approach of encoding exemplar trees that maximize representation of the full training set’s computation trees has proven to be highly effective. This strategy not only overcomes the paradoxical requirement of training on full datasets for condensation but also eliminates the need for repeated condensation when changing hyper-parameters or Gnn architectures. Furthermore, Bonsai achieves substantial size reduction without resorting to fully-connected, edge-weighted graphs, thereby reducing computational demands. In contrast to baselines, Bonsai is completely CPU-bound, resulting in at least 17-times lower carbon emissions, making it a more environmentally friendly option. These features collectively position Bonsai as a significant advancement in the field of graph condensation, offering both performance benefits and sustainability advantages.
Limitations and Future Works: While existing research on graph condensation has primarily focused on node and graph classification tasks, Gnns have demonstrated their versatility across a broader spectrum of applications. In our future work, we aim to expand the scope of graph condensation by developing task-agnostic data condensation algorithms.
@@
To support the reproducibility of our work, we provide several resources in the paper and its supplementary materials. The source code of our models and algorithms is available at https://github.com/idea-iitd/Bonsai, which also includes details of how to train the baseline models in different settings. All theoretical results and assumptions are detailed in the Appendix A, ensuring clarity in our claims. Full experimental settings are documented in Appendix B.
Acknowledgements
We acknowledge the Yardi School of AI, IIT Delhi for supporting this research. This work was partially supported by the CSE Research Acceleration Fund of IIT Delhi. Sayan Ranu acknowledges the Nick McKeown Chair position endowment. Hariprasad Kodamana acknowledges IIT Delhi Abu Dhabi for partial support of this research. Samyak Jain acknowledges the generous grant received from Microsoft Research India to sponsor his travel to ICLR 2025; and Hudson River Trading LLC for funding his work on the project.
@@ A.3 Proofs of Monotonicity, Submodularity
A.4 Sparsification through Personalized PageRank (PPR)
Let $\mathcal{V}_{S}$ be the node set in the initial condensed graph, which can be partitioned into two disjoint subsets: $\mathcal{V}_{root}$ and $\mathcal{V}_{ego}$, where $\mathcal{V}_{root}=\{v\mid\exists\mathcal{T}_{v}^{L}\in\mathbb{A}\}$, the root nodes, that serve as the root node of some exemplar in $\mathbb{A}_{greedy}$, and $\mathcal{V}_{ego}=\mathcal{V}_{S}\setminus\mathcal{V}_{root}$, the ego nodes, that are included because they fall within the $L$-hop neighborhood of a root node. We propose to employ Personalized PageRank (PPR) (Page et al., 1998) with teleportation only to root nodes to identify and prune ego nodes with minimal impact on the root node embeddings. Starting with timestamp $t=0$, we proceed as follows:
  1. 1.
    Compute the PPR distribution $\bm{\pi}_{t}$ for $t\geq 0$ defined by:
    $\bm{\pi}_{t}=(1-\beta)\mathbf{A}\bm{\pi}_{t-1}+\beta\mathbf{e}$ (27)
  2. 2.
  3. 3.
  4. 4.
    Define the knee point $i_{knee}$ as:
  5. 5.
    Prune all ego nodes with PPR scores below $\bm{\pi}[i_{knee}]$.
  6. 6.
  7. 7.
A.5 Complexity Analysis
@@
    Coverage Maximization: Each iteration in Alg. 1, iterates over all nodes (trees) in the graph ($\mathbb{T}$) that have not yet been added to the set of exemplars. For each node (equivalently, tree rooted at this node), we compute its marginal gain (line 3 in Alg. 1), which is bounded by the sample size $z$, since the cardinality of a Rev-$k$-NN set is bounded by the number of trees sampled for Rev-$k$-NN computation ((§ 3.1.1)). Since $z\ll n$, the complexity per iteration is $\mathcal{O}(zn)\approx\mathcal{O}(n)$. The number of iterations is bounded by the size of the condensation budget, which we typically set to less than $3\%$ of the full dataset and hence has negligible impact on the complexity.
@@
    PPR: PPR can be computed in linear time (Lofgren et al., 2014). Since the size of the condensed set is at most the full graph, this bounds the cost of PPT to $\mathcal{O}(n)$.